III. John is currently 25 years old. He has $10,000 saved up and wishes to deposit this into a savings account which pays him J12 = 6% p.a. He also wishes to deposit $X every month into that account so that when he retires at 55, he can withdraw $2000 every month end to support his retirement. He expects to live up till 70 years. How much should he deposit every month into his account?
1
Expert's answer
2020-06-03T19:52:12-0400

The future value is:

"FV = 15\u00d712\u00d72,000 - 10,000 = 350,000."

The monthly payment is:

"Pmt = \frac{350,000\times0.06/12}{(1 + 0.06/12)^{30\times12} - 1} = 348.43."